The price of a company's share dropped by 3.00% by the end of the first year, down to $42.25. During the second year the price of the share dropped by $1.69.

a. What was the price of the share at the beginning of the first year?

Round to the nearest cent


b. What was the price of the share at the end of the second year?

Round to the nearest cent


c. What was the percent change in the price of the share over the two years?

%



1
Expert's answer
2021-05-25T17:33:41-0400

a. Let x be the price of the share at the beginning of the first year.

Price of share at the end of the first year = x - 3.0% of x

"42.25= x -0.03x \\

42.25= 0.97x \\

x = \frac{42.25}{0.97}"

= $43.55

b. During the second year the price of the share dropped by $1.69

The price of the share at the end of the second year = $43.55 - $1.69 = $41.86

c. Price difference = $43.55 – $41.86 =$1.69

Proportion:

43.55 – 100 %

1.69 – x

"x = \frac{1.69 \times 100}{43.55}= 3.88 \;\%"

The percent change in the price of the share over the two years is 3.88 %.